A hand exerciser utilizes a coiled spring. A force of 92.9 N is required to compress the spring by 0.0201 m. Determine the force needed to compress the spring by 0.0513 m.
Answer #1

for a spring


F = k*x


F1 = k*x1


F2 = k*x2

F2/F1 = x2/x1

F2/92.9 = 0.0513/0.0201

F2 = 237.1 N